### v1

# Distributed Learning Dynamics Converging to the Core of -Matchings

Abstract
-matching is a special case of matching problems where nodes can join multiple matchings with the degree of each node constrained by an upper bound, the nodeβs -value. The core solution of a bipartite -matching is both a matching between the agents respecting the upper bound constraint and an allocation of the value of the edge among its nodes such that no group of agents can deviate and collectively gain higher allocation. We present two learning dynamics that converge to the core of the bipartite -matching problems. The first dynamics are centralized dynamics in the nature of the Hungarian method, which converge to the core in a polynomial time. The second dynamics are distributed dynamics, which converge to the core with probability one. For the distributed dynamics, a node maintains only a state consisting of (i) its aspiration levels for all of its possible matches and (ii) the matches, if any, to which it belongs. The node does not keep track of its history nor is it aware of the environment state. In each stage, a randomly activated node proposes to form a new match and changes its aspiration based on the success or failure of its proposal. At this stage, the proposing node inquires about the aspiration of the agent it wants to match with to calculate the feasibility of the match. The environment matching structure changes whenever a proposal succeeds. A state is absorbing for the distributed dynamics if and only if it is in the core of the -matching.
I INTRODUCTION
Matchings are used widely in multi-agent systems to model setups with pairwise interactions. Edges between agents can describe communication, contracts or assignments. Matchings have been studied in various forms, connection-wise: one-to-one, many-to-one, and many-to-many, and graph-wise: general and bipartite. Constraints and parameters can also be added to limit the number of connections and the weight/cost of these connections. The different forms of matchings allow us to reduce many applications to matching problems. For example, multi-agent task allocation problems can usually be modeled as one form of matching according to the objectives of the application[1]-[3]. Matching applications also include the widely known kidney exchange problem, the medical residents matching program in the United States, and ride-sharing applications [4]-[5].
Various real-life applications can be modeled as many-to-many bipartite matching problems, for instance, online ad placements [6]-[7] and firms and consultants matching [8]. In addition, the authors in [9] modeled data markets where data can be replicated with zero-cost as many-to-many bipartite matching. Furthermore, in the networks field, many problems are related to matching through duality, reduction, or extension, such as edge covering, stable set polyhedron, flow, transportation, and transshipment problems [10].
The author in [11] presented some key considerations for multi-agent systems. Among these considerations are the autonomy of the multi-agent system, the complexity of the decision-making and learning, the communication between agents, the distribution of the agents, and the systemβs security and privacy requirements. The distributed dynamics presented in this paper address a number of these considerations since the dynamics are distributed, require only simple computations from the agents per iteration, and reveal only minimal information about other agentsβ states. The setup we are considering is the bipartite -matching setup, which allows many-to-many matchings with an upper bound on the number of matches per node and at most one match between each pair of nodes. In our dynamics, an agent is randomly activated which proposes to match with another agent if both of their aspirations can be fulfilled by the value of the match. The proposing agent increases or decreases its aspiration according to the success or failure of the proposal and its current matching state. An agent cannot exceed its upper bound of matches, hence if it is already at its capacity for matches, it needs to break off an old match to join a new one.
The presented dynamics converge to the core allocation of -matching problems. The core allocation is an essential solution concept in cooperative games that guarantees that no group of agents can gain more by deviating and forming their own coalition in the game. The core allocation is proven to exist for bipartite -matchings in [12]. We prove that through our distributed dynamics and only pairwise interactions, the dynamics reach a core allocation with probability one.
Our distributed dynamics, the -matching proposals dynamics, extend and build upon the dynamics in [13]-[15]. The dynamics in [13] and [14] address the one-to-one bipartite matching setup and the many-to-one bipartite matching setup, respectively. The dynamics in [15] is applicable to general coalitional games with transferable utility, an example of which is the matching problems on a general graph. Similar work has been carried out in [16] for assignment games and in [17] for superadditive coalitional games. For -matchings, a core allocation can be reached using the dynamics in [15] if it is formulated as a coalitional game, yet it requires the calculations of the maximum -matching value (defined later) for each coalition of agents. Our current proposed dynamics bypass these calculations and use only pairwise interactions to reach the core of the -matching.
The rest of this paper is organized as follows. Section II introduces some preliminaries for our setup. Section III defines the Paths Transfers algorithm and exhibits its convergence proof. Section IV introduces the -matching Proposals algorithm. Section V discusses the -matching Proposals dynamics convergence. In Section VI, we illustrate the -matching proposals dynamics on a multi-agent task assignment setup. Finally, Section VII concludes the paper.
II PRELIMINARIES
II-A Definitions
Our setup in this paper is the weighted simple -matching presented in [10]. Henceforth, we will refer to it as the nodes -matching. The following nomenclature will be used throughout the paper.
| The grand union of all nodes | |||||
| Edges between nodes | |||||
| The -copies identifiers set | |||||
| The -th copy of node | |||||
| The grand union of all copies | |||||
| Edges between copies | |||||
| Nodes allocation | |||||
Definition 1.
Consider a complete bipartite graph together with a weighing function and a -value function Let and the degree of a node within be If for all Then is a -matching.
Definition 2.
Set the maximum -matching value of a graph and the associated and functions to be where is the set of -matchings over the graph. The maximum -matching value over a coalition where and is the maximum -matching value of over the subgraph with the associated and functions.
The following definition follows the definition of the core of -matching in [12].
Definition 3.
An allocation is a core allocation of the -matching problem if both of the following conditions are satisfied. (i) For any coalition where and the sum of the allocation of the nodes in the coalition is greater than or equal to the maximum -matching value over the coalition. (ii) The total allocation of all nodes in is equal to the maximum -matching value over the coalition, i.e. where is the set of -matchings over the complete graph.
We associate with the core allocation a -matching that achieves the maximum -matching value and denote the pair as a nodes-core solution.
As an illustration, consider a multi-agent task assignment setting, where each agent is able to perform multiple tasks and each task can be performed by several agents. The profit that is realized from an agent performing a task depends on both the agent and the nature of the task. The previous definitions can be interpreted as follows. The two node groups and are the group of agents and the group of tasks, respectively. A connected edge is an agreement that agent performs task The profit gained as agent performs task is the weight of the edge, The profit of matched edges can then be divided among the agents and the tasks. The -value for a task is an upper bound on the maximum number of agents that can perform the task and the -value for an agent is the maximum number of tasks it can perform. The allocation dictates how much gain each agent and task get from the profits collected from all matched edges.
In this paper, we are considering further a different perspective for the -matching. We represent each node as copies and specify the copies of the nodes that are connected. We define the -matching and the core from this perspective as follows.
Definition 4.
Consider a complete bipartite graph together with a weighing function and a -value Consider and as defined above. Let the degree for all and for each and then is a -copies matching.
Definition 5.
A reduction function is a function that maps a pair of copies allocation and -copies matching to a pair of nodes allocation and -matching as follows. If then and Note that is still a -matching as previously defined because for each node
From this paperβs perspective, each node will be expanded to distinct copies, each treated as a node in an expanded graph and gets its own gain from the profit of the matches; a nodeβs gain then is the sum of the gains collected by all of its copies.
Definition 6.
A copies-core solution is a pair of copies allocations and -copies matching that satisfies the following three conditions.
1.
Edge saturation:
2.
Stability against pairwise deviation:
Let then -
3.
Zero-gain for unmatched copies:
Proposition 1.
If is a copies-core solution then is a nodes-core solution.
Proof.
1.
Consider any be a subset of and be the edges over
2.
Let
β’
be a max nodes -matching for the set i.e. where is a -matching over
β’
be the value of such -matching where and
β’
β’
3.
Let and
4.
Construct in the following way.
β’
Loop over For each pick any free and i.e. neither matched in nor in and add to
β’
5.
β’
For all since
β’
For all since then and
β’
6.
This implies that
7.
Due to edge saturation and zero-gain for unmatched copies, the following holds.
Since is a -copies matching, then,
From point
where is a maximum -matching over
Finally, since is a maximum -matching and the total aspirations of the copies equal to a -matching, then from the previous inequality and equation, we get thatwhere is a maximum -matching over
From points and and the definition of the reduction function, then
and
Hence, is a nodes-core solution.
Consider the copies-core solution of the previous task assignment setting. For this solution, an allocation is specified for each copy of the node. This finer detailing of the copiesβ allocations shows the gains of each node, robot or task, from each match it is part of, thus, decoupling the individual matchesβ profit distribution. In a copies-core solution, the allocation specified for each copy of a node is only its profit allocation from the match that this copy is a part of.
In the next two sections, we define a centralized polynomial time algorithm as well as distributed dynamics for finding the copies-core solution. Since we are aiming to reach the copies-core solution, we work with the expanded graph as defined above. The following definitions will be used in the subsequent algorithms. Furthermore, we will overload in the dynamics to be the agentsβ allocation aspirations for the copies. In our dynamics, an edge is formed only if these aspirations can be fulfilled by such match.
Definition 7.
The free copies set in a matching is defined as Node free copies are Additionally, denote the set of free copies with strictly positive aspirations as and for any node
Definition 8.
The partner function in a matching is defined as where and if and if
The following is an assumption that we are going to consider for both the centralized and distributed algorithm.
Assumption 1.
For all , and for all meaning that the nodes cannot have more copies than the number of nodes in the opposite class.
If a given -matching problem does not satisfy this assumption, then the degree of the violating nodes still will not exceed the number of nodes in the opposite class in any -matching, and hence, the problem remains the same if we decrease the -values of these nodes just enough to satisfy the above assumption.
The algorithms use the aforementioned functions, however, the superscript are suppressed and considered as the matching at the current iteration.
III PATHS TRANSFERS ALGORITHM
In this section, we define the paths transfers algorithm, which is a centralized algorithm that is in the nature of the Hungarian method [10] and influenced by BLMAβs convergence proof in [13]. The algorithm converges in a polynomial time to a copies-core solution. The algorithm proceeds as follows.
1.
Initially, the matching will be empty and the aspirations of all copies will be . Set to be the largest discretization value for all the weights of the edges.
2.
Over-aspiration step: In this stage, the algorithm loops over all nodes in one nodesβ class, without loss of generality, we will set that to be the class. For each set the aspiration of all copies to the maximum weight of the edges that has as an endpoint, i.e. set for all and
3.
Outer loop: Loop while
β’
Pick any without loss of generality assume and denote this copy by If the same subsequent steps should be followed exchanging the and instances with the and instances.
β’
Inner loop:
Loop so long as has not decreased. Consider the current matching let and construct the equality directed graph as follows:
β
if
β
if
Check the following cases in order. If one occurs, follow the steps of that case.
i.
Decreasing aspiration case: If there is no edge in directed from to any then decrease by
If reaches zero, then decreased by 1, then continue with the outer loop of step Otherwise, reconstruct and repeat step inner loop considering the same
ii.
Augmenting path case: If there exists a directed path in from the considered copy to some copy then let and disregarding the directions of the edges. Add edges in to and remove from
This case will lead to decreasing by 1 or 2 depending whether is greater than or equal to
iii.
Copies exchange case: If there exists a directed path in from the considered copy to some copy with then let and disregarding the directions of the edges. Add edges in to and remove from
This case will lead to decreasing by 1 as a copy of zero aspiration will be exchanged with a copy of non-zero aspiration in the matched copies.
iv.
Aspiration transfer case: If none of the above cases occur, then start with and construct the following sets. Let and If and are non-empty, then let and Keep on constructing and as long as is non-empty, as follows. Let and Let the final constructed sets be and Decrease the aspiration of all copies in in addition to by and increase the aspiration of all copies in by
One implementation of the above steps will decrease the total aspirations of copies in and increase the total aspirations of copies in The step must end once no more copies can be added to or which must happen since old copies in previous sets cannot be added to This step may not lead to a decrease in if so, repeat step βs inner loop starting from the same and reconstruct .
β
The inner loop of step must end in a finite time. The inner loop will end if it goes through the augmenting path or copies exchange cases or if the considered copy reaches zero aspiration in the decreasing aspiration case. The inner loop cannot infinitely go through the decreasing aspiration and the aspiration transfer case since the total aspirations of copies in strictly decreases with each implementation of these cases and it is bounded below by zero. Hence, in a finite number of steps, the inner loop will end up in the considered copy reaching zero aspiration and hence decreases or the loop goes through the augmenting path or copies exchange cases.
β’
The outer loop of step must end because strictly decreases with each iteration.
Theorem 1.
The Paths Transfers algorithm reaches the copies-core solution in a finite number of steps.
Proof.
First, the initialization and over-aspiration step guarantees that we reach a state that satisfies both the edge saturation and stability against pairwise deviation conditions. The algorithm then aims to satisfy the zero-gain for unmatched copies condition while keeping the first two conditions satisfied. After each iteration of the outer loop of step both edge saturation and stability against pairwise deviation conditions remain satisfied. Consider one implementation of the outer loop of step .
β’
For the decreasing aspiration case, no change in the matching occurs, and the considered copy decreases its aspiration since it strictly satisfies the stability against pairwise deviation. Since we decrease by on an -discritized grid, the copy will remain satisfying the stability against pairwise deviation, and since it is not matched, it will not affect the edge saturation condition.
β’
For the augmenting path and copies exchange case, new connections are only made between copies whose total aspirations are exactly equal to the weight of the edge between them, since these edges belong to the equality graph. Hence, the edge saturation condition is preserved. Note that all copies on the paths in these cases have minimum aspirations compared to the aspirations of the other copies of their respective nodes. In addition, the summation of aspirations of these, now unmatched, copies equal to the weight of the edge between them. Thus, these copies along with other copies of these two nodes should still preserve the stability against pairwise deviation condition.
β’
For the aspiration transfer case, the matching does not change, and for each edge considered that is both in the matching and the equality graph, one endpoint decreases its aspiration by epsilon while the other endpoint increases its aspiration by epsilon, hence keeping their total aspiration equal to the weight of the edge and preserving the edge saturation condition. In addition, for each copy that decreases its aspiration, any copy of the opposite class that may break the stability against pairwise deviation condition, tightly satisfies the condition with before the decrease in its aspiration, increases its aspiration by hence, preserving this condition as well.
Finally, since the outer loop of step in the algorithm strictly decreases then the algorithm will eventually stop and the zero-gain for unmatched copies condition will be satisfied. Hence, the final state will be a copies-core solution.
Note that the algorithm can be modified to be polynomial in and and exclude the weight values dependency. To achieve that, the aspiration transfer case should be modified such that the copies aspirations change by the minimum change of aspiration that leads to a new edge in the equality graph to appear or for one of the copies to hit zero aspiration. The decreasing aspiration case should be modified such that the decrease in aspiration be the minimum decrease such that a reconstructed contains an edge from the considered copy to any other copy or its aspiration reaches zero.
IV -MATCHING PROPOSALS ALGORITHM
In this section, we introduce the -matching proposals dynamics. Intuitively, the algorithm follows these steps
1.
Agents come with arbitrary non-negative initial aspirations for all their copies. The aspirations should lie on a grid of width , a discretization value for all the weights of the edges.
2.
The algorithm then iterates over the following steps:
i.
An agent is activated uniformly at random, which in turn chooses, using a uniform distribution, an agent in the opposite class of nodes to propose forming a match with.
ii.
If the current agents are matched, the proposal is skipped, otherwise, the proposing agent asks for and receives, an aspiration value from the other agent.
iii.
The receiving agent sends the minimum aspiration of its unmatched copies, if any exist, otherwise, it sends the minimum aspiration of the matched copies. Similarly, the proposing agent considers the minimum aspiration of its unmatched copies, if any exist, otherwise, the minimum aspiration of the matched copies. If the total of the received aspiration in addition to the proposing agentβs considered aspiration raised by is less than or equal to the proposed match weight, a match between the two copies of the considered aspirations is formed, otherwise, the proposal fails.
a.
If a match is successfully formed, the proposing agent increases its aspiration to be the difference between the weight of the edge between the proposer and the receiver and the receiverβs aspiration. If any of the copies in the new match were already matched, it dissolves its previous match to join the new one.
b.
If a match proposal fails, the proposing agent decreases its aspiration in one of the copies that are not matched and has an aspiration greater than zero, if any exist, by
a.
i.
V -MATCHING PROPOSALS ALGORITHM ANALYSIS
Definition 9.
A state is feasible if is a -copies matching, for all and for all
Proposition 2.
Starting from a feasible state and following the -matching proposals algorithm, the state remains feasible.
Proof.
At each iteration, one proposal is made, if the proposal succeeds, then the aspiration of the copies forming the new matching exactly equals the weight of the edge between them. In addition, the aspirations of the agents remain on the -grid. If the proposal fails, only agents with strictly positive aspiration decrease their aspiration by and since the aspirations are on the -grid, the aspiration will either be greater than or equal to zero. Furthermore, if there exist copies of a pair of nodes that are already matched, no further copies for the same nodes can match, and each copy is an endpoint of at most one edge in the match, hence, the -copies matching and feasibility conditions are preserved.
Theorem 2.
The -matching Proposals algorithm converges to the copies-core solution and hence, by reduction, to the nodes-core solution with probability one.
Proof.
The proof follows three steps illustrated in the following subsections. First, the there exists a positive probability bounded away from zero to reach a copies-core solution from any feasible state within some finite time . Second, the algorithm converges to a copies-core solution with probability one. Third, a state is absorbing if and only if it is a copies-core solution.
I. There exists a positive probability to reach a copies-core solution from any feasible state
Proposition 3.
Starting from any feasible state, there exists a finite sequence of proposals resulting in a state satisfying the edge saturation and stability against pairwise deviation conditions.
Proof.
Each node can propose with a positive probability to another node of the opposite class that it is not currently matched with. Consider the sequences of proposals, where nodes (from one class) propose to form an edge only whenever there is a possible gain in the proposing agentβs aspiration. Such sequences will have a finite number of proposals due to the boundedness of the number of edges, and hence, will occur with positive probability bounded away from zero within a time horizon equal to the number of edges, . Any of these sequences of proposals will result in a state satisfying edge saturation, due to the statesβ feasibility, and stability against pairwise deviation conditions, since these sequences only stop when there are no more proposals that can result in aspiration gain.
Proposition 4.
Starting from any feasible state satisfying the edge saturation and stability against pairwise deviation conditions, there exists a finite sequence of proposals resulting in a copies-core solution.
Proof.
The proof follows from the fact that there exists a sequence of proposals that occurs with positive probability, within a finite time horizon that is polynomial to the input parameters. Such possible sequence corresponds to the implementation of step in the paths transfers algorithm. This sequence consists of subsequences that are analogous to performing the outer loop of step and lead to strictly decreasing after each subsequence. At the beginning of each subsequence, similar to the paths transfer algorithm, the equality graph, and the cases that the state can be in, guide the possible subsequence of proposals that leads to a decrease in .
The first part of each subsequence is analogous to the decreasing aspiration case as follows. For each node with free copies that have aspirations greater than zero, at least one of these copies should be part of the equality graph. If that is not the case for one node, then the proposal sequence in this part is as follows. Without loss of generality assume such node belongs to and denote it by then consider such that then let propose to repeatedly until either such that and or a copy of hits zero aspiration from the failed proposals leading to a strict decrease in Recalculate the equality graph, and repeat for any other node with the same violation, exchanging the and instances with and if such node belongs to instead of If decreases after the first part, then the subsequence ends then. Otherwise, the subsequence has a second part that corresponds to one of the following three cases.
β’
If there exists a directed path on the equality graph from a free copy from one class to a copy of the same class with zero aspiration, then a set of proposals analogous to the copies exchange case is followed. The proposals are such each node on this path proposes to the subsequent node on the path twice, starting from the free copy tailing this path and excluding the proposal from the second to last node to the last node. The subsequence ends after this set of proposals since it leads to strictly decreasing.
β’
If such directed path does not exist but an augmenting path exists from a free copy with positive aspirations, then the case becomes analogous to the augmenting path case. A similar set of proposals is followed where each node on this path proposes to the subsequent node on the path twice, starting from the free copy tailing this path and including the second to last node proposing to the last node. The subsequence ends after this set of proposals since it leads to strictly decreasing.
β’
If none of these paths exist, then a set of proposals analogous to the operation in the aspiration transfer case is followed. The following set of proposals leads to the total aspirations of one class of nodes decreasing and the total aspirations of the other class increasing. One set of such proposals does not guarantee a decrease in yet repeating the set with the same class of nodes as the starting point would eventually either lead to decreasing and the subsequence ends then or to a directed path on the updated equality graph to appear following one of the cases that were mentioned earlier. If such directed path appears, the corresponding set of proposals can be followed to decrease and end the subsequence. The aforementioned set of proposals that is analogous to the operation in the aspiration transfer case proceeds as follows.
β
This set of proposals is guided by choosing a node with a free copy of positive aspiration from one class of nodes. Without loss of generality, consider such node a node denoted by and and are constructed in the same manner as in the Paths Transfers algorithm.
β
The first proposal in this set is the proposal of to a node such that and The proposal fails because of the satisfaction of the stability against pairwise condition. Hence, the aspiration of a free copy, with minimum aspiration, of decreases. Now let be the node neighbors of through the equality graph, i.e. all the nodes that have copies as heads for edges that have a copy of as the tail. Call the following set of proposals a U proposal round. Let each node in propose to in order. The proposals should succeed, old matches get broken, and the aspiration for a copy of each proposing node increases. After this round of proposals each that had a match broken due to these proposals tries to restore that match, denote this round by the matches restoration round. In this round, nodes in try to restore broken matches by proposing back to their old partners. The proposals initially fail and the total aspiration of the proposing nodes decreases. Another two proposals are made in an attempt to restore these matches, whenever possible by either the node from or whichever node has free copies. The U proposal and match restoration rounds are repeated as needed till and either stay the same or decreases compared to its value at the beginning of the subsequence.
β
Repeat the previous U proposal and match restoration rounds but for each of instead of If the considered had a match before beginning the rounds that are pertaining to it, then after finishing these rounds, the node that was connected to proposes twice to it to restore that old match after the end of these rounds.
β
II Borel-Cantelli lemma
Proposition 4 implies that there exists time for which the probability of reaching a copies-core solution from any feasible state is at least some that does not depend on the state, following the aforementioned finite sequences. Hence, by the Borel-Cantelli lemma, the algorithm reaches a copies-core solution with a probability one.
III. A state is absorbing if and only if it is a copies-core solution
If a state is absorbing then it is a copies-core solution
Assume the state is absorbing. Let For a state to be absorbing then,
β’
the edge saturation property is satisfied since it is satisfied for all feasible states.
β’
and the zero-gain for unmatched copies condition is satisfied. Otherwise, such that and due to Assumption there exists a node in the opposite class that is not connected to through any copy. If proposes to match with such node, the proposal should fail since the state is absorbing and the matching structure cannot change, yet one copy of from will decrease its aspiration, hence the state is not absorbing, which provides a contradiction.
β’
Given , then and the stability against pairwise deviation is satisfied. Otherwise, and such that thus if proposes to or vice versa, the proposal will succeed changing the matching and the aspirations. Hence the state is not absorbing, which provides a contradiction.
Consequently, is a copies-core solution.
If a state is a copies-core solution then it is absorbing
Assume the state is a copies-core solution at iteration . Assume that at iteration any was chosen to propose to any
β’
If and were already connected: no change occurs.
β’
If and are not connected, then by the stability against pairwise deviation property, and Hence, all proposals will fail and no change can occur in the matching .
In addition, from zero-gain for unmatched copies condition, , then none of the copies can decrease its aspiration.
Similar justification is applied if any proposes to any Hence, is absorbing.
VI ILLUSTRATION
In this section, we demonstrate the distributed learning dynamics on the aforementioned multi-agent task assignment setting. The setup is as follows.
β’
We consider tasks, where
β
each task has a value such that and
β
can afford to utilize up to robots
β
β’
We consider robots, where
β
each robot has accuracy value where and
β
can afford to cover up to tasks.
β
in our setup, we are assuming a negative correlation with the sophistication of the robot and how many tasks it can cover
β
β’
The value of a matching of task with robot is where is proportional with the accuracy of the robot multiplied by the value of the task
VI-A Sample run and average performance
The illustrations shown are algorithm runs with a setup of the following main parameters:
β’
tasks with ranging from to
β’
robots with ranging from to
β’
The left side of Fig. 1 shows the final matching between the robots nodes, with each node expanded as copies, and the tasks nodes, with each node expanded as copies. The right side of Fig. 1 has a plot of the total aspirations of all the copies involved in a matching at each iteration. The dashed line represents the maximum -matching value solved by a linear program.
In Fig. 2, we ran the algorithm for different configurations, with the same parameters as above but with randomization of the values and -values of the tasks. The figure shows the plot of the average of the relative total feasible aspirations, where, for each run, the relative total feasible aspiration is the total feasible aspirations divided by the optimal value for that run.
VI-B Parameters effects
We ran various simulations to help us gain insights into the effects of some of the problem parameters on the average performance of the dynamics, specifically the -values, number of nodes, and discretization size. Figures 3, 4, and 5 illustrate the observed trends resulting from the parameters variations.
First, to isolate the effect of varying the -value, we generated random -matching instances, and for each instance, we ran the dynamics for different -value functions, and . value is randomly generated such that is such that and is such that . Fig. 3 shows the average of the relative welfare across runs while using the different -value functions. Empirically, it appears for this setup that bigger -values lead to better performance of the dynamics, on average.
Second, we generated random -matching instances, and for each instance, we ran the dynamics for different discretization values. Fig. shows the average of the relative welfare across runs while using the different discretization values. From the simulations, we observed that the smaller the discretization value is, the longer the dynamics took to converge and get closer to the welfare value.
Finally, we varied the number of nodes for random -matching instances to observe the impact of decreasing the number of nodes on the dynamics performance. For each instance, we removed a number of nodes gradually and their associated edges, reran the dynamics, and compared the empirical performance. Fig. 5 illustrates the different performance levels for the dynamics resulting from varying the number of nodes.
VII CONCLUSION
To recap, we considered the -matching setup. We defined a copies core solution, which is a solution concept that specifies the gain of each node in a bipartite graph, from each match it joins. The copies-core solution achieves core stability such that no group of agents can deviate and collectively gain a higher payoff through another matching. We presented a centralized algorithm, in the nature of the Hungarian method, that reaches the copies-core solution in polynomial time. We further introduced distributed learning dynamics that converge to the core with probability one. Finally, we illustrated the distributed dynamics on a multi-robot task assignment setting.